OpenStudy (anonymous):

Solve for x: 4|2x − 2| + 10 = 34

OpenStudy (anonymous):

a. x = −4, x = 2 b. x = 4, x = −4 c. x = 2, x = −2 d. x = 4, x = −2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

absolute value there is no negatives

OpenStudy (anonymous):

@kyle55

OpenStudy (anonymous):

could i ask you a few more

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

Consider the functions f(x) = 3x2 + 1 and g(x) = x + 3. What is the value of f[g(−1)]? Select one: a. 7 b. 37 c. 13 d. 12

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

@_greatmath7 please explain how you got that answer

OpenStudy (texaschic101):

4 | 2x - 2 | + 10 = 34 --- subtract 10 4 | 2x - 2 | = 24 -- divide by 4 | 2x - 2 | = 6 2x - 2 = 6 -(2x - 2) = 6 2x = 6 + 2 -2x + 2 = 6 2x = 8 -2x = 6 - 2 x = 4 -2x = 4 x = -2

OpenStudy (anonymous):

and for the second one?

OpenStudy (texaschic101):

sorry...not good with functions

OpenStudy (welshfella):

you can find f(g(x) first - by replaceing the x in f(x) by g(x)

OpenStudy (welshfella):